How do I simplify this and express in scientific notation:
The ratio 35,960,000 miles : 3,664,000,000 miles, the ratio of the distance of the sum from Mercury to its distance from Pluto
Help!

Answer by checkley75(3666)

35,960,000=3.596*10^7 THIS SAYS THE DECIMAL POINT WAS MOVED 7 PLACES TO THE LEFT.
3,664,000,000=3.664*10^9 MOVED THE DECIMAL 6 PLACES TO THE LEFT.
-----------------------------------------------------------------------
now we divide the numbers & subtract the exponents (7-9)=2. the -9 is due to it being in the denominator.
3.596*10^7/3.664*10^9
.9814*10^-2 or
9.814*10^-3 the answer.